OpenStudy (anonymous):

solve please 2-2(3a+6) = 4a-(a-2)

OpenStudy (anonymous):

2-2(3a+6) = 4a-(a-2) 2-6a-12 = 4a-a+2 -6a-4a+a = 2-2+12 9a = 12 a = 4/3 or 1.33

OpenStudy (anonymous):

my answer booklet says a= -4/3

OpenStudy (anonymous):

-6a-4a+a = 2-2+12 -9a = 12 => a = -12/9 = -4/3 There you are!

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Sorry for the signs.. It should be -4/3!